OpenStudy (anonymous):

Find the angle from the first line to the second. 3x = 4 , 3x + 2y = -8

OpenStudy (anonymous):

can you help me please i will give medal :D

OpenStudy (owlcoffee):

The angle between two lines is defined by: \[Cos \alpha =\frac{ \left| \vec{u}. \vec{v} \right| }{ \left| \vec{u} \right|\left| \vec{v} \right| }\] Quite the thing huh? but u and v, represent the direction vectors of the line in question, the first ti the second, that's the order. But let's rewrite that equation because it is a little messy that way (I will do it in such way that i will be easier to replace), so let's see the following: \[\left| \vec{a}. \vec{b} \right|=(x _{a}x _{b})+ (y _{a}y _{b})\] And we can calculate the module of a vector, in a ortonormated reference point as: \[\left| \vec{b} \right|=\sqrt{x _{b }^{2}+y _{b}^{2}}\] So, replacing: \[Cos \alpha = \frac{ (x_u x_v + y_u y_v ) }{ \sqrt{x _{u}^{2} +y _{u}^{2} }\sqrt{x _{v}^{2}+y _{v}^{2}} }\]

OpenStudy (owlcoffee):

So, let's find the directional vectors of those lines, let's call them "u" and "v": We can find the directional vector of any line by the following equation: \[ax +by +c=0\] \[ \vec{w}(-b,a) \] so, we can get the directional vector with the (-b,a) form, you can also use (b,-a) but I personally prefer the first one. So, taking the lines: \[3x=4 \rightarrow 3x-4=0\] so therefore, the vector "u" is: \[ \vec{u}(0,3)\] And the second line: \[3x + 2y=8 \rightarrow 3x +2y-8=0\] The second vector "v" is: \[ \vec{v}(-2,3)\] Now, it's just a matter of applying the equation I stated above.

OpenStudy (anonymous):

i think its ( 2, -3)

OpenStudy (owlcoffee):

what form do you use to find the directional vecotr of a line?

OpenStudy (anonymous):

A slope

OpenStudy (anonymous):

(0,3) , (2, -3) isnt it?

OpenStudy (owlcoffee):

Sure, we can go with that. Try placing it in the equation I wrote in the first post.

OpenStudy (anonymous):

i will use the slope formula to get the angle :D

OpenStudy (anonymous):

m = y2-y1/ x2-x1

OpenStudy (owlcoffee):

Okay, let me see your work when you're finished.

OpenStudy (anonymous):

then m = tan data :D

OpenStudy (anonymous):

i got a negative angle

OpenStudy (anonymous):

-71.57 degrees

OpenStudy (anonymous):

is this right?

OpenStudy (owlcoffee):

Well, I believe you're using: \[m=\tan \alpha\] But that alpha there, is the angle the line has referential to the x-axis. Like this: 

OpenStudy (owlcoffee):

so what you calculated is that little angle, you did well, but it's not the angle yo uare asked for.

OpenStudy (anonymous):

how can we get it ? D:

OpenStudy (owlcoffee):

I gave you the formula, in the first post I made.

OpenStudy (anonymous):

ok, i will solve for it

OpenStudy (anonymous):

i got 33.68 degrees

OpenStudy (owlcoffee):

If you did the maths correctly, that should be correct.

OpenStudy (anonymous):

:D , can you help me another run

OpenStudy (owlcoffee):

sure!